# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### 26 Jan 2024 v3

@@ 4.2 Risk-coping Mechanisms
Preceding results show that the estimated income elasticities show that the skilled workers’ households in Tsukishima were less likely to smooth their consumption. The estimated elasticities, however, become smaller when I focus on the semi-monthly frequency. Specifically, food expenditure is likely to be insured within at least half a month. The workers might have been able to deal with their shorter-run shocks relatively well but could not cope with the longer-run shocks that bring lasting (clear) impacts on their income in at least one month. In this section, I test the mechanisms behind their smoothing behavior using several variables on the risk-coping devices.
@@
where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household fixed effect, $\zeta_{t}$ indicates the time fixed effect, and $\epsilon_{i,t}$ is a random error term. Online Appendix C.1 summarizes the derivation of this specification. I prefer to use the head’s income instead of the disposable income in equation 2 to avoid the potential endogeneity. [^61] For instance, the household’s total income is the outcome of the intra-household responses to the shocks, such as the labor-supply adjustments. I confirmed that the results are largely unchanged if I use the disposable income as the idiosyncratic income shock variable instead of the head’s income (not reported). This is consistent with the fact that the labor-supply adjustments did not occur among the THBS households (Table 5). Despite this, I use the conservative setting to deal with potential biases. See Asdrubali et al. (2020) for a recent study using the head’s income as the shock variable in testing the full risk-sharing. As for $r_{i,t}$, I consider several variables to overview the wide variety of risk-coping strategies. These include the net income on savings, insurance, borrowing, credit purchases, and gifts. Each net variable is defined as the difference between income and expenses. The estimates of $\kappa$ are expected to be negative since net income becomes negative in good times but is positive if the breadwinner’s earnings drop. I also consider the sales of miscellaneous goods and the earnings of all the family members except for the head because asset sales and intra-household labor supply adjustments may play a role in mitigating the idiosyncratic shocks as alternative risk-coping strategies (Udry 1995; Blandell et al. 2016). The estimated coefficients for the sales and labor supply adjustment variables are expected to be negative, reflecting additional earnings when the head’s income is reduced. Panel B of Table 3 lists the summary statistics of these variables.
@@ Results
(Each net variable is defined as the expenditure minus income)
Net savings -0.036 [0.016]** 300 -0.028 [0.016]* 174
Net insurance 0.000 [0.014] 300 0.017 [0.024] 174
Net borrowing -0.021 [0.026] 300 0.034 [0.037] 174
Net credit purchase -0.108 [0.049]** 300 -0.012 [0.060] 174
Net gifts -0.052 [0.019]*** 300 -0.013 [0.015] 174
@@
Other member’s earnings 0.001 [0.010] 300 0.037 [0.024] 174
Sales of assets -0.000 [0.000] 300 0.000 [0.000] 174
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: This table shows the results of equation 2: the family size control, household fixed effects, and time-fixed effects are included in all the regressions. Columns 1 and 2 show the estimates for the semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on the head’s earnings are listed in the “Coef.” columns. Panel A: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit redeemed minus the credit expenses in the credit purchase. “Net gifts” refer to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts. Panel B: “Other family member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
@@
The estimated coefficient for the net savings is negative and statistically significant in Column (1). The estimate suggests that a one yen decrease (increase) in the head’s earnings increases (decreases) the net income from savings by $0.036$ yen. The estimate from the monthly series is slightly smaller ($0.028$) but still weakly statistically significant (Column (2)). This is consistent with the historical fact that many workers had savings accounts both in postal savings and savings banks at that time (Section 2.3). However, the estimated magnitudes are considerably small, suggesting that the withdrawals from savings did not provide sufficient amounts of temporary income to compensate for the losses of earnings.
@@
The estimates for the insurance category are close to zero and statistically insignificant in both columns. The budget books indicate that the primary source of in category is the mutual loan associations (mujin), a local savings institution among workers. As explained, these associations’ rules were unsuitable for compensating the idiosyncratic shocks (Section 2.2). A suggestive case is that a THBS household won a lottery and got $100$ yen in a monthly cell, clearly independent of the worker’s earnings.626262 The OISR, Archives of the Tsukishima Survey (THBS #5, 17th May 1919). The bidding rule may be used to support a specific member in the association if members have a common belief in the aid. However, most of the incomes from mujin measured in the budget books are less than one yen, which is not the case on average.
@@
For the net borrowing category, the estimated coefficients are negative but statistically insignificant in both columns. Section 2.2 shows that pawnshop was the popular lending institution, especially for lower-income working-class households. Although THBS households occasionally used pawnshops, they did not use them frequently. [^63] The OISR, Archives of the Tsukishima Survey (THBS #2; 19; 22). This statement is logical because THBS households are not categorized as belonging to the lower-income working class.
@@
The estimated coefficient for the net income from credit purchases is negative and weakly statistically significant in Column (1). The estimate suggests that a one yen decrease (increase) in the head’s earnings increases (decreases) the net income from credit purchases by $0.1$ yen. This magnitude is substantially larger among the other categories. According to the THBS budget books, the credit purchases were mostly used to buy rice, fish, vegetables, and fuels (firewood and charcoal). Since rice, a staple food, is usually an expensive item, credit purchases might have played an important role among these households. While this provides a suggestive evidence that credit purchase is the primary risk-coping strategy among these households at that time (Section 2.2), Column (2) shows that credit purchases were less likely to work in the longer-run. This seems to be reasonable because borrowers could make monthly or semi-monthly payments but could not borrow for more than a month. Instead, the result for the net savings category suggests that precautionary savings have helped mitigate income declines lasting at least one month due to idiosyncratic shocks. Given the magnitudes, however, it is important to note that both devices could only partially compensate for the idiosyncratic shocks.
@@
The estimates for the net income from gifts are negative and statistically significant in Column (1). The THBS budget books show that the gifts included seasonal gifts, celebration money, and return gifts. Among them, the return gifts can be regarded as mutual aid from personal networks, which may be reflected in this result. According to an official report of Tokyo, however, the share of average monthly gifts in total income was approximately $3$% (Social Affairs Division 1925, pp. 58-59). In my adjusted monthly panels, this share is also calculated to be roughly $4$%. In this light, while the gifts comprised a few percentage points of the monthly income, they may not have provided sufficient aid for the households facing economic hardship. In fact, the estimate is no longer statistically significant in the adjusted monthly series (Column (2)).
@@
Finally, Panel B tests the roles of the intra-household labor supply adjustment and sales of miscellaneous assets. The estimated coefficients in both columns are close to zero and statistically insignificant, indicating no labor supply adjustments occurred. Although sewing was the most popular side job among THBS wives, their earnings were too small to replace the lost income from their breadwinning husbands. [^64] The OISR, Archives of the Tsukishima Survey (THBS #1, 13th January 1919). In addition, I find that the sales of miscellaneous assets such as newspapers, bottles, and clocks were not used for the risk-coping.
@@
To summarize, I find suggestive evidence that credit purchases have played a role in mitigating the idiosyncratic shocks. The results also indicate that the savings institutions and personal networks might have been used as risk-coping institutions among the workers. However, these institutions were basically available for short-term idiosyncratic shocks lasting at most half a month, albeit the savings institutions may be used for dealing with the longer shocks. The estimated magnitudes indicate that the workers could only partially compensate for the temporary losses of income by using these institutions, which is indeed consistent with the findings on the partial risk-sharing in Section 4.1. Overall, the results are aligned with the descriptives from the historical documents on the financial institutions available in Tokyo City (Section 2.3). An interesting finding, which is not explicitly described in the historical sources, is the possibility that the informal gifts from personal networks might have provided some aid to the factory workers.
@@ 4.3 Event-Study Analysis: Response to the Advers Health Shocks
The foregoing analyses used the within variations in the income variables for estimating the income elasticity, meaning that the identification depends both on the adviser and favorable idiosyncratic shocks. While it is useful to illustrate the overall responses of the households to the idiosyncratic shocks, it does not offer a clear view of the pathways behind the risk-coping behaviors against the adverse idiosyncratic shocks. To clarify this point, I next focus on the temporary illness of the heads as the adverse idiosyncratic shocks that cause temporary income losses.
@@
Notes: This figure shows the time-series plots of the monthly death counts from influenza (ryūkosei kanbō) in Tokyo prefecture between January 1918 and December 1920. The death statistics used in this figure were collected based on the comprehensive registration system. Source: Created by the author using the 1918, 1919, and 1920 editions of the Statistics of Causes of Deaths of the Empire Japan (Statistics Bureau of the Cabinet 1921; 1922; 1923).
@@
I begin my analysis by testing whether the illness of the head reduced the earnings. To do so, I decided the exact timing and duration of illnesses of the family members using the Kazoku idō hyō (Section 3.4). Figure 5 shows the number of monthly death counts from influenza (ryūkōsei kanbō) in Tokyo between 1918 and 1920. There appears to have been a moderate spike in early 1919, which is partially captured in the THBS period. In fact, a few THBS budget books suggest that the family members had caught the flu during this period. Five households experienced the illness during the sample period, and the average duration for the heads is $5.8$ days (Std. Dev. $=3.6$). Type A influenza causes fever for less than 6 days (Uyeki et al. 2022), and the illness affecting the heads was only reported between February and May 1919, when the exposure was most prevalent. [^65] Among four cells with health shocks on the heads, two are observed in February, during the peak of the epidemics (Figure 5), and the others are in March and May. Therefore, it is plausible to see that the influenza epidemics largely led the illnesses during this period. Considering this feature, this historical event offers me a practical quasi-experimental setting in analyzing to what extent the adverse idiosyncratic shock driven by illness influences the households’ consumption behaviors.
@@
Table 6 presents the results from the regressions of the head’s earnings on the number of days suffering illness in the previous semi-month cell. One must be careful that the head’s earnings in a specific semi-month cell depend on the works in the last semi-month cell. In other words, if the health shocks have reduced their earnings, the lagged health shock alone has a clear effect on the earnings, whereas the instantaneous health shock does not. Columns (1)–(3) confirms this mechanism. While the instantaneous health shock has no impact on the head’s earnings (Column 1), the lagged health shock has a statistically significant effect on his earnings (Column 2). This result is unchanged if I include both variables in the same regression in Column 3, supporting the evidence that the health shock reduced the earnings of the head. The estimate in Column (2) suggests that a one-day increase in the illness decreases the head’s semi-monthly income by roughly $2.5$ yen. Notably, this is close to the average daily wage rate of skilled factory workers (Online Appendix B.6). Thus, this may reflect the reductions in their salary due to the absence. Finally, Column (4) adds the health shocks of the other family members into the specification of Column (3). The estimated coefficients on these placebo variables are statistically insignificant, supporting the robustness of the result in Column (3). These results confirm that the head who suffered illness lost earnings. [^66] Given that the average number of days suffering the illness is roughly $2.0$, the temporary health shock loses more than $5$ yen in that month, accounting for roughly $10$% of the average monthly income. Although this is a rough estimate and the losses shall depend on the severity of the shocks, my result at least shows that the household income had reacted to the adverse health shock among the factory workers’ households.
@@
  The head [level] $0.064$ $-0.739$ $-0.823$
$[0.585]$ $[0.668]$ $[0.679]$
  The head [lag] 2.449***-2.449***- 2.449 * * * 2.827***-2.827***- 2.827 * * * 2.876***-2.876***- 2.876 * * *
$[0.586]$ $[0.670]$ $[0.684]$
  The others [level] $0.195$
$[0.642]$
  The others [lag] $0.418$
$[0.655]$
Observations 298 263 263 263
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: This table shows the results from the regressions of the head’s earnings on the health shock variables using the semi-monthly panel dataset. Since the lagged health shock variables are used, a household with two semi-month observations is excluded from the sample in all the regressions. Therefore, the number of THBS households used in this analysis is 33. The family size control, household fixed effects, and time-fixed effects are included in all the regressions. “Health shocks on the head” refers to the number of days that the head suffered illness. ”Health shocks on the others” refers to the number of days that family members, excluding the head, were ill. “[level]” indicates the variable of the head’s earnings in the same semi-month cell. “[lag]” indicates the lagged variable (i.e., the shocks measured in the previous semi-month cell).
Following the evidence, I characterize equation 1 for the consumption smoothing regression as a structural equation by using the first-stage reduced form equation as follows:
@@
where $h_{i,t-1}$ is the number of days of head’s illness in the previous semi-month cell, $\varpi_{i}$ is the household fixed effect, $\tau_{t}$ is the semi-month fixed effect, and $e_{i,t}$ is a random error term. For the risk-sharing mechanisms analysis under equation 2, the log-transformed disposable income in equation 3 is replaced with the head’s earnings. The health shock variable is plausibly exogenous because the illnesses were from the influenza epidemics, as explained above. In addition, there is a correlation between the health shock variable and the head’s earnings at a statistically valid level (Table 6). Both support the evidence that the exogeneity of the instrument and relevance condition basically hold under this just-identified system. Panel D of Table 3 summarizes the summary statistics of the health shock variables.
@@
Total consumption -0.117 [0.015]*** 263 0.534 [0.068]*** 253 18.9
    Food -0.075 [0.017]*** 245 0.347 [0.072]*** 235 20.0
    Housing -0.017 [0.070] 129 0.116 [0.469] 129 7.8
    Utilities -0.052 [0.067] 207 0.199 [0.330] 199 12.3
    Furniture -0.149 [0.099] 164 0.715 [0.742] 157 6.6
    Clothes -0.339 [0.050]*** 227 1.233 [0.222]*** 218 65.4
    Education -0.099 [0.039]** 185 0.371 [0.152]** 178 46.2
    Medical expenses -0.001 [0.028] 245 -0.010 [0.144] 235 20.0
    Entertainment expenses 0.027 [0.045] 232 -0.239 [0.460] 222 15.2
    Transportation -0.100 [0.094] 177 0.436 [0.507] 169 21.3
    Gift -0.323 [0.091]*** 219 1.559 [0.784]** 210 18.4
    Miscellaneous -0.077 [0.026]*** 204 0.348 [0.100]*** 197 18.8
@@
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: Column 1 shows the results of equation 1, in which the disposable income ($y_{i,t}$) is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable. The dependent variables are the 11 measures of log-transformed consumption, and all the regressions include the family size control, household fixed effects, and time fixed effects. The estimated coefficients on the health shock variable are listed in the column named “Coef.”. Column 2 shows the results of equation 1, in which the disposable income ($y_{i,t}$) is regarded as the endogeneous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) for the first-stage reduced-form equation. The estimated coefficients on the disposable income are listed in the column named “Coef.”. The first-stage $F$-statistics on this health shock variable are reported in the final column.
@@
I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1. Column 1 in Table 7 shows the results. Overall, the coefficients on the exposure variable are negative, supporting the evidence that adverse health shocks on the heads reduce consumption levels. The estimate for the food subcategories is smaller as I found in my baseline result (Table 4). The estimate for the clothing category shows a relatively large magnitude, which is consistent with the adjustments in reducing luxury consumption. The expenses on gifts were also likely to be cut off in response to the adverse health shock as those are optional. Interestingly, the estimate for the education category is now statistically significantly negative. This may reflect the reduction of allowance for children in response to the income losses due to the health shocks.

### 25 Feb 2024 v6

@@ 4.2 Risk-coping Mechanisms
Preceding results show that the skilled workers’ households in Tsukishima were less likely to smooth their consumption. The estimated elasticities, however, become smaller when I focus on the semi-monthly frequency. Specifically, food expenditure is likely to be insured within at least half a month. The workers might have dealt with their shorter-run shocks relatively well but could not cope with the longer-run shocks. In this section, I test the mechanisms behind their smoothing behavior using several variables on the risk-coping devices.
@@
where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ is the household fixed effect, $\zeta_{t}$ is the time fixed effect, and $\epsilon_{i,t}$ is a random error term. Online Appendix C.1 summarizes the derivation of this specification. I prefer to use the head’s income instead of the disposable income in equation 2 to avoid the potential endogeneity. [^60] For instance, the household’s total income is the outcome of the intra-household responses to the shocks, such as the labor-supply adjustments. I confirmed that the results are largely unchanged if I use the disposable income as the idiosyncratic income shock variable instead of the head’s income (not reported). This is consistent with the fact that the labor-supply adjustments did not occur among the THBS households (Table 5). Despite this, I prefer to use the conservative setting to deal with potential biases and to set a comparable setting to the structural equation used in Section 4.3. See Asdrubali et al. (2020) for a recent study using the head’s income as the shock variable in testing the full risk-sharing. As for $r_{i,t}$, I consider several variables to overview various risk-coping strategies. These include the net income on savings, insurance, borrowing, credit purchases, and gifts. Each net variable is defined as the difference between income and expenses. The estimates of $\kappa$ are therefore expected to be negative since net income becomes negative in good times but positive in unfavorable times. I also consider the earnings of all the family members except for the head and the sales of miscellaneous goods because both intra-household labor supply adjustments and asset sales may play a role in mitigating the idiosyncratic shocks as alternative risk-coping strategies (Udry 1995; Blundell et al. 2016). [^61] Rosenzweig and Stark (1989) show that marriage cum migration in South Indian villages contributed to reducing the variability of household food consumption. Since the THBS focuses on the households living in Tsukishima throughout the sample period, I can discard addressing the ex-ante risk-coping strategy through migration. The estimated coefficients for the sales and labor supply adjustment variables are expected to be negative, reflecting additional earnings when the head’s income is reduced.
@@ Results
Net savings -0.043 [0.018]** 289 -0.042 [0.018]* 162
Net insurance 0.001 [0.013] 289 0.015 [0.024] 162
Net borrowing -0.021 [0.027] 289 0.036 [0.036] 162
Net credit purchase -0.128 [0.039]*** 289 -0.017 [0.059] 162
Net gifts -0.052 [0.019]** 289 -0.015 [0.014] 162
@@
Other member’s earnings 0.001 [0.010] 289 0.038 [0.024] 162
Sales of assets -0.000 [0.000] 289 0.000 [0.000] 162
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: This table shows the results of equation 2: the family size control, household fixed effects, and time-fixed effects are included in all the regressions. Columns 1 and 2 show the estimates for the semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on the head’s earnings are listed in the “Coef.” columns. Panel A: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts. Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
@@
The estimated coefficient for the net savings is negative and statistically significant in Column (1). The estimate suggests that a one yen decrease (increase) in the head’s earnings increases (decreases) the net income from savings by $0.043$ yen. The estimate from the monthly series is similar ($0.042$), albeit weakly statistically significant (Column (2)). This is basically consistent with the historical fact that many workers had savings accounts both in postal savings and savings banks at that time (Section 2.3). However, the estimated magnitude is small, suggesting that the withdrawals from savings did not compensate for the earnings losses.
@@
The estimates for the insurance category are close to zero and statistically insignificant in both columns. The budget books indicate that the primary source in the category is the mutual loan associations (mujin). As explained in Section 2.2, these associations’ rules were unsuitable for compensating the idiosyncratic shocks. A suggestive case is that a THBS household won a lottery and got $100$ yen in a monthly cell, which is clearly independent of the worker’s earnings. [^62] The OISR, Archives of the Tsukishima Survey (THBS #5, 17th May 1919). The bidding rule may be used to support a specific member in the association if members have a common belief in the aid. However, most of the incomes from mujin measured in the budget books are less than one yen, which is not the case on average.
@@
For the net borrowing category, the estimated coefficients are close to zero and statistically insignificant in both columns. Section 2.2 shows that pawnshop was the popular lending institution for low-income households. Although THBS households occasionally used pawnshops, they never used them frequently. [^63] The OISR, Archives of the Tsukishima Survey (THBS #2; 19; 22). This makes sense because skilled factory workers can not classified as the low-income working class.
@@
The estimated coefficient for the net income from credit purchases is negative and statistically significant in Column (1). The estimate suggests that a one yen decrease (increase) in the head’s earnings increases (decreases) the net income from credit purchases by $0.13$ yen. This magnitude is larger among the other categories. According to the THBS budget books, the credit purchases were mostly used to buy rice, fish, vegetables, and fuels (firewood and charcoal). Since rice, a staple food, is usually an expensive item, credit purchases might have played an important role among these households. The credit purchases were occasionally used in the call doctor and pharmacy. This is consistent with my baseline result for medical expenses in the semi-monthly series (Column (1) in Table 4). While this provides suggestive evidence that credit purchase is the primary risk-coping strategy, Column (2) shows that the credit was less likely to work in the longer run. This seems to be reasonable because borrowers could make monthly or semi-monthly payments but did not borrow for more than a month (Section 5). Instead, the result for the net savings category suggests that precautionary savings have helped mitigate income declines lasting at least one month due to idiosyncratic shocks. Given the magnitudes, however, it is important to note that it could only partially compensate for the idiosyncratic shocks.
@@
The estimates for the net income from gifts are negative and statistically significant in Column (1). The budget books tell that the gifts included seasonal gifts, celebration money, and return gifts. Among them, the return gifts can be regarded as mutual aid from personal networks, which may be reflected in this result. According to an official report of Tokyo, however, the share of average monthly gifts in total income was approximately $3$% (Social Affairs Division 1925, pp. 58-59). In my adjusted monthly panels, this share is calculated to be roughly $4$%. In this light, while the gifts comprised a few percentage points of the monthly income, they may not have provided sufficient aid for the households facing economic hardship. In fact, the estimate is no longer statistically significant in the adjusted monthly series (Column (2)).
@@
Finally, Panel B tests the roles of the intra-household labor supply adjustment and sales of miscellaneous assets. The estimated coefficients in both columns are close to zero and statistically insignificant. Although sewing was the most popular side job among THBS wives, their earnings were too small to replace the lost income from their breadwinning husbands. [^64] The OISR, Archives of the Tsukishima Survey (THBS #1, 13th January 1919). In addition, I find that the sales of miscellaneous assets such as newspapers, bottles, and clocks did not work as the risk-coping.
@@
To summarize, I find that credit purchases have played a role in mitigating the idiosyncratic shocks. The results also indicate that the savings institutions and personal networks might have been used as risk-coping institutions among the workers. However, these were basically available for short-term, say at most half a month. The estimated magnitudes indicate that the workers could only partially compensate for the temporary losses of income by using these institutions, which is indeed consistent with the findings on the partial risk-sharing in Section 4.1. Overall, the results are aligned with the descriptives from the historical documents on the financial institutions available in Tokyo City (Section 2.3). An interesting finding, which is not explicitly described in the historical sources, is the possibility that the credit purchases in the local society and the informal gifts from personal networks might have provided some aid to the factory workers.
@@ 4.3 Event-Study Analysis: Response to the Advers Health Shocks
The foregoing analyses used the within variations in the earnings for estimating the income elasticity, meaning that the identification depends both on the adverse and favorable idiosyncratic shocks. While it is useful to illustrate the overall responses of the households to the shocks, it does not offer a clear view of the pathways behind the risk-coping behaviors against the adverse idiosyncratic shocks. To clarify this point, I next focus on the temporary illness of the heads that cause temporary income losses.
@@
Notes: This figure shows the time-series plots of the monthly death counts from influenza (ryūkōsei kanbō) in Tokyo prefecture between January 1918 and December 1920. The death statistics used in this figure were collected based on the comprehensive registration system. Source: Created by the author using the 1918, 1919, and 1920 editions of the Statistics of Causes of Deaths of the Empire Japan (Statistics Bureau of the Cabinet 1921; 1922; 1923).
@@
I begin my analysis by testing whether the illness of the head impacted the earnings. To do so, I decided the exact timing and duration of illnesses of the family members using the Kazoku idō hyō (Section 3.4). Figure 5 shows the number of monthly death counts from influenza (ryūkōsei kanbō) in Tokyo between 1918 and 1920. There appears to have been a moderate spike in early 1919, which is partially captured in the THBS period. In fact, a few THBS budget books suggest that the family members had caught the flu during this period. Five households experienced the illness during the sample period, and the illness affecting the heads was concentrated between February and May 1919, when flu exposure was most prevalent. [^65] Among four cells with health shocks on the heads, two are observed in February, during the peak of the epidemics (Figure 5), and the others are in March and May. In addition, the average duration for the heads is $5.8$ days (Std. Dev. $=3.6$), which is consistent with the fact that Type A influenza causes fever for less than $6$ days (Uyeki et al. 2022). Therefore, it is plausible to see that the influenza epidemics largely led to illnesses during this period, providing a quasi-experimental setting in analyzing to what extent the adverse idiosyncratic shock driven by illness influences the households’ consumption behaviors.
@@
Earnings in a semi-month are based on the work done during the previous semi-month. This means that the head’s earnings shall mainly be influenced by the number of days suffering illness in the previous semi-month cell. Table 6 presents the results from the regression of the head’s earnings on the number of illness days in the previous semi-monthly cell. While the instantaneous health shock has little impact on the head’s earnings (Column 1), the lagged health shock has a statistically significant effect on his earnings (Column 2). This result is unchanged if I include both variables in the same regression in Column 3, supporting the evidence that the health shocks in the previous cell reduced the earnings of the head. The estimate in Column (2) suggests that a one-day increase in the illness decreases the head’s semi-monthly income by roughly $2.4$ yen. Notably, this is close to the average daily wage rate of skilled factory workers (Online Appendix B.6). Thus, this may reflect the reductions in their salary due to the absence. Finally, Column (4) adds the health shocks of the other family members into the specification of Column (3). The estimated coefficients on these placebo variables are statistically insignificant, supporting the robustness of the result in Column (3). These results confirm that the head who suffered illness lost earnings. [^66] Given that the average number of days suffering the illness is roughly $2.0$, the temporary health shock loses more than $5$ yen in that month, accounting for roughly $10$% of the average monthly income. Although this is a rough estimate and the losses shall depend on the severity of the shocks, my result at least shows that the household income had reacted to the adverse health shock among the factory workers’ households.
@@
  The head [level] $-0.061$ $-0.972$ 1.319*-1.319*- 1.319 *
$[0.679]$ $[0.699]$ $[0.666]$
  The head [lag] 2.371***-2.371***- 2.371 * * * 2.868***-2.868***- 2.868 * * * 2.994***-2.994***- 2.994 * * *
$[0.644]$ $[0.734]$ $[0.765]$
  The others [level] $0.889$
$[1.275]$
  The others [lag] $3.675$
$[3.312]$
Observations 287 254 254 254
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: This table shows the results from the regressions of the head’s earnings on the health shock variables using the semi-monthly panel dataset. A household with two semi-month observations is excluded from the sample in all the regressions because the lagged health shock variables are included. The family size variable, household fixed effects, and time-fixed effects are included in all the regressions. “Health shocks on the head” refers to the number of days that the head suffered illness. “Health shocks on the others” refers to the number of days that family members, excluding the head, were ill. “[level]” indicates the variable of the head’s earnings in the same semi-month cell. “[lag]” indicates the lagged variable (i.e., the shocks measured in the previous semi-month cell). Although correlation coefficient between the level and lag of the health shock variables for the other family members is substantially high (Pearson’s coefficient $=0.99$), including eather of the two variables does not change the finding on the neglisible impacts of the health shocks on the other family members.
Following the evidence, I characterize equation 1 as a structural consumption smoothing equation by using the first-stage reduced form equation as follows:
@@
where $h_{i,t-1}$ is the number of days of head’s illness in the previous semi-month cell, $\varpi_{i}$ is the household fixed effect, $\tau_{t}$ is the semi-month fixed effect, and $e_{i,t}$ is a random error term. For the risk-sharing mechanisms analysis under equation 2, the log-transformed disposable income in equation 3 ($y_{i,t}$) is replaced with the head’s earnings ($\tilde{y}_{i,t}$). The health shock variable is plausibly exogenous because the illnesses were from the influenza epidemics, as explained above. In addition, there is a correlation between the health shock variable and the head’s earnings at a statistically valid level (Table 6). Both support the evidence that the exogeneity of the instrument and relevance condition basically hold under this just-identified system.
@@
Total consumption -0.118 [0.015]*** 254 0.543 [0.076]*** 245 17.6
    Food -0.075 [0.017]*** 236 0.355 [0.075]*** 227 18.4
    Housing -0.017 [0.070] 126 0.118 [0.477] 126 7.6
    Utilities -0.051 [0.066] 199 0.201 [0.331] 192 12.4
    Furniture -0.157 [0.106] 156 0.761 [0.801] 150 6.2
    Clothes -0.339 [0.050]*** 219 1.233 [0.222]*** 211 65.4
    Education -0.099 [0.039]** 176 0.373 [0.153]** 170 45.9
    Medical expenses -0.002 [0.028] 236 -0.001 [0.142] 227 18.4
    Entertainment expenses 0.019 [0.046] 224 -0.222 [0.501] 215 13.1
    Transportation -0.100 [0.094] 170 0.436 [0.510] 162 20.5
    Gift -0.325 [0.093]*** 212 1.567 [0.797]* 204 17.8
    Miscellaneous -0.075 [0.027]*** 195 0.343 [0.101]*** 189 17.3
@@
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: Column 1 shows the results of equation 1, in which the log-transformed disposable income is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable. The dependent variables are the 11 measures of log-transformed consumption, and all the regressions include the family size control, household fixed effects, and time fixed effects. The estimated coefficients on the health shock variable are listed in the sub-column named “Coef.”. Column 2 shows the results of equation 1, in which the log-transformed disposable income is regarded as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) for the first-stage reduced-form equation (3). The estimated coefficients on the disposable income are listed in the sub-column named “Coef.”. The first-stage $F$-statistics on the health shock variable are listed in the final column.
@@
I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1 in the spirit of Cochrane (1991). Column 1 in Table 7 shows the results. Overall, the coefficients on the exposure variable are negative, supporting the evidence that adverse health shocks on the heads reduce consumption levels. The estimate for the food subcategory is relatively small. The estimates for the clothing and gifts categories show a relatively large magnitude. These are in line with my baseline results in Table 4. The estimate for the education category is now statistically significantly negative. This may reflect the reduction of allowances for children in response to income losses due to health shocks.